# Retrieval Augmented Generation for Dynamic Graph Modeling

### v1

@@
Abstract
Dynamic graph modeling is crucial for analyzing evolving patterns in various applications. Existing approaches often integrate graph neural networks with temporal modules or redefine dynamic graph modeling as a generative sequence task. However, these methods typically rely on isolated historical contexts of the target nodes from a narrow perspective, neglecting occurrences of similar patterns or relevant cases associated with other nodes. In this work, we introduce the Retrieval-Augmented Generation for Dynamic Graph Modeling (RAG4DyG) framework, which leverages guidance from contextually and temporally analogous examples to broaden the perspective of each node. This approach presents two critical challenges: (1) How to identify and retrieve high-quality demonstrations that are contextually and temporally analogous to dynamic graph samples? (2) How can these demonstrations be effectively integrated to improve dynamic graph modeling? To address these challenges, we propose RAG4DyG, which enriches the understanding of historical contexts by retrieving and learning from contextually and temporally pertinent demonstrations. Specifically, we employ a time- and context-aware contrastive learning module to identify and retrieve relevant cases for each query sequence. Moreover, we design a graph fusion strategy to integrate the retrieved cases, thereby augmenting the inherent historical contexts for improved prediction. Extensive experiments on real-world datasets across different domains demonstrate the effectiveness of RAG4DyG for dynamic graph modeling.
1 Introduction
@@
Figure 1: Illustration of RAG in NLP and dynamic graph modeling. (a) In NLP, RAG leverages pre-trained language models to encode text and retrieve semantically similar or related demonstrations, which are further concatenated to enhance the generation task. (b) Our work addresses the challenges of complex temporal and structural characteristics of dynamic graphs, incorporating RAG through time- and context-aware retrieval and graph fusion modules.
Dynamic graphs, which track the temporal evolution of nodes and edges, are essential for analyzing complex systems across domains like social networks, citation networks etc. Existing methods for dynamic graph modeling fall into two main categories: discrete-time and continuous-time approaches (Feng et al. 2024). Discrete-time methods divide the timeline into intervals, capturing the graph’s state at each point (Sankar et al. 2020), but they lack temporal granularity by overlooking event dynamics within intervals. Continuous-time approaches, on the other hand, model temporal dynamics continuously, without discretization, allowing for more detailed representations (Trivedi et al. 2019; Xu et al. 2020; Wen and Fang 2022). Our work aligns with continuous-time approaches, aiming to capture fine-grained time steps for more accurate dynamic graph modeling.
Traditional approaches for dynamic graph modeling predominantly rely on integrating graph neural networks (GNNs) with specialized temporal modules, such as recurrent neural networks (Pareja et al. 2020), self-attention (Sankar et al. 2020) and temporal point processes (Wen and Fang 2022), which update the graph representations over time. Although powerful, these methods suffer from the inherent limitations of GNNs, such as difficulty with long-term dependencies, and over-smoothing or over-squashing issues (Chen et al. 2020; Alon and Yahav 2020). Recently, a simple and effective approach called SimpleDyG (Wu, Fang, and Liao 2024) redefines dynamic graph modeling as a generative sequence modeling task and leverages the strengths of sequence models like Transformers to capture long-range dependencies within temporal sequences effectively.
Despite significant advancements in dynamic graph modeling, current approaches heavily depend on isolated historical contexts of the target nodes, limiting each node’s perspective to its ego network. The limited context makes it difficult to quickly adapt to diverse emerging patterns. For example, in a social network, when users transit to new behavioral patterns or when new users with minimal historical interactions are introduced, traditional models that focus on the historical contexts of individual nodes might not respond effectively to these changes. To overcome these limitations, we draw inspiration from the Retrieval-Augmented Generation (RAG) technique originated from the Natural Language Processing (NLP) field (Gao et al. 2023). As shown in Fig. 1(a), RAG broadens the context through the retrieval of additional demonstrations, achieving considerable success in NLP. In this work, we aim to integrate RAG into dynamic graph modeling to incorporate a broader and more relevant contextual understanding beyond the historical interactions of individual nodes. However, leveraging RAG for dynamic graph modeling presents two critical challenges.
First, how do we harvest high-quality demonstrations for dynamic graphs? This involves selecting the most contextually and temporally relevant samples to enrich the current state of a target node. Unlike RAG for NLP, where external textual data and a pre-trained language model (LM) naturally support demonstration retrieval as shown in Fig. 1(a), dynamic graph modeling requires careful consideration of the data sources and the ability to handle complex structural and temporal patterns as shown in Fig. 1(b). To address this, we introduce a novel time-aware contrastive learning strategy that leverages internal training data for demonstration retrieval. Following SimpleDyG (Wu, Fang, and Liao 2024), we regard dynamic graph modeling as a sequence modeling task and the prediction of future events as a sequence generation problem. Hence, we define a query sequence for a target node as its historical interaction sequence. Next, we automatically annotate the retrieval training pool by assessing the similarity to the query. Given the training pool, our model consists of two contrastive modules: time-aware and context-aware contrastive learning. The former utilizes a time decay function to prioritize samples that are temporally close to the query, while the latter incorporates data augmentation techniques such as masking and cropping to improve the model’s ability to retrieve complex structural patterns.
Second, how do we effectively integrate the retrieved demonstrations into the dynamic graph model? This involves fusing the retrieved demonstrations to enrich the historical context of each target node and enhance subsequent predictions. Directly concatenating the retrieved demonstrations with the query sequence not only results in a lengthy sequence unsuited for generative predictions but also overlooks the structural patterns among the retrieved demonstrations. To address this, we propose graph fusion, leveraging the graph structures inherent to the demonstrations and fusing them into a summary graph, as shown in Fig. 1(b). We then apply a GNN-based readout to learn the representation of the fused summary graph, which is prepended to the query sequence before being fed into the sequence generation model.
@@
Our contributions can be summarized as follows. (1) We propose a novel Retrieval-Augmented Generation approach for Dynamic Graph modeling named RAG4DyG. It employs a retriever to broaden the historical interactions with contextually and temporally relevant demonstrations, leading to enhanced generative predictions on dynamic graphs. (2) We propose a time-aware contrastive learning module that incorporates temporal and structural information for demonstration retrieval. Additionally, we design a graph fusion module to effectively integrate the retrieved demonstrations, leveraging structural enhancements for subsequent predictions. (3) We conduct extensive experiments to validate our approach, demonstrating the effectiveness of RAG4DyG in various domains.
@@
Figure 2: Overall framework of RAG4DyG. (a) Sequence modeling for dynamic graphs. (b) The retriever finds top-K temporally and contextually relevant demonstrations. (c) Graph fusion integrates the retrieved demonstrations for subsequent generation.
2 Related Work
Dynamic Graph Modeling.
@@ 2 Related Work
Existing approaches for dynamic graphs can be categorized into discrete-time and continuous-time methods. Discrete-time methods regard a dynamic graph as a sequence of static graph snapshots captured at various time steps. Each snapshot represents the graph structure at a specific time step. These methods typically adopt GNNs to model the structural information of each snapshot, and then incorporate a sequence model (Pareja et al. 2020; Sankar et al. 2020) to capture the changes across snapshots. However, these approaches neglect fine-grained time information within a snapshot. In contrast, continuous-time methods model graph evolution as a continuous process, capturing all time steps for more precise temporal modeling. These methods often integrate GNNs with specially designed temporal modules, such as temporal random walk (Wang et al. 2021), temporal graph attention (Xu et al. 2020; Rossi et al. 2020), MLP-mixer (Cong et al. 2022) and temporal point processes (Trivedi et al. 2019; Ji et al. 2021; Wen and Fang 2022). Recently, researchers have proposed a simple and effective architecture called SimpleDyG (Wu, Fang, and Liao 2024), which reformulates the dynamic graph modeling as a sequence modeling task. Specifically, it maps the dynamic graph into a series of node sequences and feeds them into a generative sequence model. Subsequently, predicting future events can be framed as a sequence generation problem.
Retrieval Augmented Generation.
@@
Recently, the RAG paradigm has attracted increasing attention (Gao et al. 2023). Specifically, RAG first leverages the retriever to search and extract relevant documents from some databases, which then serve as additional context to enhance the generation process. Related studies have demonstrated the great potential of RAG in various tasks such as language processing (Karpukhin et al. 2020; Jiang et al. 2023), recommendation systems (Ye et al. 2022; Dao et al. 2024), and computer vision (Liu et al. 2023; Kim et al. 2024).
In the graph modeling field, existing RAG efforts have primarily focused on static and text-attributed graphs to enhance the generation capabilities of Large Language Models, supporting graph-related tasks such as code summarization (Liu et al. 2021) and textual graph question answering (He et al. 2024; Hu et al. 2024). However, exploiting RAG techniques for dynamic graphs and graphs without textual information remains largely unexplored.
3 Preliminaries
@@
Sequence Mapping of Dynamic Graphs.
We denote a dynamic graph as $G=(V,E,F,\mathcal{T})$ comprising a set of nodes $V$, edges $E$, a node feature matrix $F$ if available, and a time domain $\mathcal{T}$. To map a dynamic graph into sequences, we follow SimpleDyG (Wu, Fang, and Liao 2024). Specifically, let $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ denote the set of training samples, where each sample is a pair $(x_{i},y_{i})$, representing the input and output sequences for a target node $v_{i}\in V$. The input $x_{i}$ is a chronologically ordered sequence of nodes that have historically interacted with $v_{i}$, while the output $y_{i}$ is the ground truth interactions that occurs following the sequence $x_{i}$. In notations, we have
@@
Problem Formulation.
Dynamic graph modeling aims to learn a model that can predict the future interactions of a target node $v_{i}$, given its historical interactions. That is, given $x_{i}$ in Eq. (1), the task is to predict $y_{i}$ in Eq. (2).
In our RAG framework, we regard the training samples $D$ as a retrieval pool. Given a target node $v_{q}\in V$, its input sequence $x_{q}$ is referred to as the query sequence. We first retrieve $K$ demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for each query sequence $x_{q}$ based on their contextual and temporal relevance. Next, the retrieved demonstrations $R_{q}$ are used to enrich the input sequence $x_{q}$, which encompasses the historical interactions of the target node $v_{q}$. The augmented input $\{R_{q},x_{q}\}$ is designed to enhance the predictions of future events in $y_{q}$.
4 Proposed Model: RAG4DyG
Our proposed RAG4DyG framework, depicted in Fig. 2, first trains a generative sequence model for dynamic graphs in Fig. 2(a). We use SimpleDyG (Wu, Fang, and Liao 2024) as the backbone for this task. As described in Sect. 3, a dynamic graph $G$ is represented as sequences of node interaction records, which are then used to train a Transformer-based sequence model.
Next, given a query sequence $x_{q}$ and a retrieval pool $D$, a time- and context-aware retriever is designed to retrieve demonstrations for $x_{q}$ from $D$, as shown in Fig. 2(b). The retriever fine-tunes its encoder (i.e., the sequence model) to jointly optimize two contrastive losses: a time-aware loss that employs a time decay function, and a context-aware loss that employs sequence augmentation. (Sect. 4.1)
Finally, given top-K demonstrations retrieved from the retrieval pool, we design a graph fusion module as illustrated in Fig. 2(c). Specifically, we fuse the K demonstrations into a summary graph, leveraging the graph structures inherent to the retrieved demonstrations. The summary graph is further encoded by a GNN, serving as an augmented context that is prepended to the original query sequence. The augmented sequence is subsequently input to the sequence model to predict future events. (Sect. 4.2)
4.1 Time- and Context-Aware Retriever
@@
Retrieval Annotation.
To facilitate contrastive training, we automatically annotate the samples in the retrieval pool $D$. For each query sequence $x_{q}$, we annotate its positive sample $x_{p}^{+}$ from the pool $D$ based on their contextual similarity. Specifically, we adopt the sequence model pre-trained in Fig. 2(a) as the encoder and apply mean pooling to obtain sequence representations. Given a query sequence $x_{q}$ and a candidate sequence $x_{p}\in D$, we define their contextual similarity as the dot product of their representations:
$s(x_{q},x_{p})=f(x_{q})^{\top}f(x_{p}),$ (3)
where $f(\cdot)$ denotes our encoder. We leave the detailed annotation process in Appendix A.
Time-aware Contrastive Learning.
Temporal information reflects the dynamic changes in historical interactions, which is crucial for dynamic graph modeling. We posit that demonstrations closer in time to the query are more relevant than those further away. Consequently, we utilize a time decay function to account for temporal proximity between the query and candidate sequences, as follows.
$\mu(t_{q},t_{p})=\exp(-\lambda|t_{q}-t_{p}|),$ (4)
@@
To effectively capture the temporal dynamics of the graph, we incorporates temporal proximity to reweigh the contextual similarity in the contrastive loss:
$h(x_{q},x_{p})$ $=s(x_{q},x_{p})\mu(t_{q},t_{p}).$ (5)
@@
$\mathcal{L}_{\text{tcl}}$ $=-\log\frac{\exp(h(x_{q},x_{p}^{+}))/\tau}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp(h(x_{q},x_{j}))/\tau},$ (6)
@@
Context-aware Contrastive Learning.
To better capture the inherent contextual pattern, we further adopt context-aware contrastive learning with data augmentations. For each sequence, we apply two types of augmentations: masking and cropping, which are widely used for sequence modeling (Devlin et al. 2019; Xie et al. 2022). The masking operator randomly replaces a portion of the tokens in the sequence with a special masking token. The cropping operator randomly deletes a contiguous subsequence from the original sequence, reducing the sequence length while preserving the temporal order of the interactions. These augmentations help the model learn robust representations and capture the inherent structural information of the sequence by focusing on its different parts.
@@
$\mathcal{L}_{\text{ccl}}=-\log\frac{\exp(s(x^{\prime}_{q},x^{\prime\prime}_{q})/\tau)}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp s(x^{\prime}_{q},x^{\prime}_{j})/\tau},$ (7)
@@
Training and Inference for Retrieval.
The training objective of our retrieval model is defined as:
$\mathcal{L}_{\text{ret}}=\mathcal{L}_{\text{tcl}}+\alpha\mathcal{L}_{\text{ccl}}$ (8)
where $\alpha$ is a coefficient that controls the balance between the two losses.
During testing, we utilize the updated sequence model to extract sequence representations and perform demonstration ranking based on the contextual similarity between the query and candidates.
4.2 Graph Fusion-based Generator
After retrieval, we obtain the top-K demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for the query $x_{q}$. A straightforward approach is concatenating them with the query sequence. However, this can lead to a lengthy context that limits the model’s prediction capabilities. More importantly, it neglects the structural patterns among these demonstrations. Thus, we first fuse the demonstrations into a summary graph, process it using a GNN, and then prepend the graph readout from the GNN to the query for subsequent generation.
Graph Fusion.

### v2

@@
Abstract.
Modeling dynamic graphs, such as those found in social networks, recommendation systems, and e-commerce platforms, is crucial for capturing evolving relationships and delivering relevant insights over time. Traditional approaches primarily rely on graph neural networks with temporal components or sequence generation models, which often focus narrowly on the historical context of target nodes. This limitation restricts the ability to adapt to new and emerging patterns in dynamic graphs. To address this challenge, we propose a novel framework, Retrieval-Augmented Generation for Dynamic Graph modeling (RAG4DyG), which enhances dynamic graph predictions by incorporating contextually and temporally relevant examples from broader graph structures. Our approach includes a time- and context-aware contrastive learning module to identify high-quality demonstrations and a graph fusion strategy to effectively integrate these examples with historical contexts. The proposed framework is designed to be effective in both transductive and inductive scenarios, ensuring adaptability to previously unseen nodes and evolving graph structures. Extensive experiments across multiple real-world datasets demonstrate the effectiveness of RAG4DyG in improving predictive accuracy and adaptability for dynamic graph modeling. The code and datasets are publicly available at https://github.com/YuxiaWu/RAG4DyG.
Retrieval-augmented generation, graph neural networks, dynamic graph modeling
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18,2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729958isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Figure 1. Illustration of RAG in NLP and dynamic graph modeling. (a) In NLP, RAG leverages pre-trained language models to encode text and retrieve semantically similar or related demonstrations, which are further concatenated to enhance the generation task. (b) Our work addresses the challenges of complex temporal and structural characteristics of dynamic graphs, incorporating RAG through time- and context-aware retrieval and graph fusion modules.
Dynamic graph modeling is essential for understanding and predicting evolving interactions across various applications such as social networks (Deng et al., 2019; Sun et al., 2022), personalized recommendations(Song et al., 2019; Wu et al., 2022; Tang et al., 2023), and Web-based services (Han et al., 2021; Duan et al., 2024). These applications require effective handling of the temporal and contextual dynamics inherent in evolving graph structures to provide accurate and timely insights. For example, in a recommendation system (Tang et al., 2023), capturing user behavioral shifts over time can enhance the personalization of content, while in social networks (Sun et al., 2022), understanding interaction patterns can improve engagement and fraud detection mechanisms.
Existing dynamic graph modeling methods generally fall into discrete-time and continuous-time approaches (Feng et al., 2024). Discrete-time models capture graph snapshots at specific intervals, providing a simplified representation of the graph’s evolution but often failing to account for fine-grained temporal dynamics (Sankar et al., 2020; Pareja et al., 2020). Continuous-time models, such as DyRep (Trivedi et al., 2019), TGAT (Xu et al., 2020), and TREND (Wen and Fang, 2022), model events as they occur, offering a more granular perspective and better capturing event-driven dynamics. These models commonly rely on graph neural networks (GNNs) integrated with temporal mechanisms such as recurrent neural networks (Pareja et al., 2020), self-attention (Sankar et al., 2020), and temporal point processes (Wen and Fang, 2022) to update graph representations over time. Despite their strengths, GNN-based methods struggle with long-term dependencies and issues such as over-smoothing and over-squashing (Chen et al., 2020; Alon and Yahav, 2020). Recently, SimpleDyG (Wu et al., 2024) redefined dynamic graph modeling as a generative sequence modeling task, leveraging Transformers to effectively capture long-range dependencies within temporal sequences. However, the reliance on localized historical interactions still limits the ability to generalize across different contexts and adapt to emerging patterns.
To address these limitations, the Retrieval-Augmented Generation (RAG) framework from the Natural Language Processing (NLP) domain (Gao et al., 2023) offers a promising approach. RAG has the potential to broaden the contextual understanding of dynamic graphs by retrieving and incorporating relevant examples from across the graph’s temporal and contextual space, as illustrated in Figure 1. However, adopting RAG for dynamic graph modeling presents two major challenges: (1) Selecting high-quality demonstrations, and (2) effectively integrating the retrieved demonstrations. Identifying contextually and temporally relevant demonstrations is challenging because existing retrieval methods, such as BM25 and other matching-based schemes (Robertson and Jones, 1976) primarily rely on historical interaction similarities and struggle with inductive scenarios where nodes lack historical interactions. Moreover, effective integration of retrieved demonstrations is another challenge, as simply concatenating them with the query sequence can lead to overly lengthy inputs and overlook underlying structural patterns.
To overcome these challenges, we propose a novel Retrieval-Augmented Generation for Dynamic Graph modeling (RAG4DyG) framework. It integrates a time- and context-aware contrastive learning strategy that evaluates historical interaction sequences to identify relevant examples and a graph fusion module to construct a summary graph from the retrieved samples. The contrastive learning strategy incorporates a time decay function to prioritize temporally relevant samples, while context-aware augmentation techniques such as masking and cropping enhance the model’s ability to capture complex structural patterns. The graph fusion module applies a GNN-based readout mechanism to enrich the representation before feeding it into the sequence generation model. These solutions empower RAG4DyG to effectively leverage retrieved demonstrations to enhance dynamic graph modeling. Through extensive experimentation on various real-world datasets, we demonstrate that RAG4DyG outperforms state-of-the-art methods in both transductive and inductive scenarios, offering improved accuracy and adaptability in dynamic graph scenarios. In transductive settings, where test nodes have appeared during training, our model effectively leverages historical data to refine predictions. In inductive settings, involving previously unseen nodes, the retrieval mechanism enables the model to generalize by providing relevant contextual examples as guidance.
@@
To sum up, our main contributions are as follows. (1) We propose a novel retrieval-augmented generation approach for dynamic graph modeling named RAG4DyG, which employs a retriever to broaden historical interactions with contextually and temporally relevant demonstrations. (2) We introduce a time- and context-aware contrastive learning module that incorporates temporal and structural information for demonstration retrieval and a graph fusion module to effectively integrate retrieved demonstrations. (3) We conduct extensive experiments to validate our approach, demonstrating the effectiveness of RAG4DyG across various domains.
@@
Figure 2. Overall framework of RAG4DyG. (a) Sequence modeling for dynamic graphs. (b) The retriever finds top-K temporally and contextually relevant demonstrations. (c) Graph fusion integrates the retrieved demonstrations for the subsequent generation.
2. Related Work
@@
Dynamic Graph Modeling. Existing approaches for dynamic graphs can be categorized into discrete-time and continuous-time methods. Discrete-time methods regard a dynamic graph as a sequence of static graph snapshots captured at various time steps. Each snapshot represents the graph structure at a specific time step. These methods typically adopt GNNs to model the structural information of each snapshot, and then incorporate a sequence model (Pareja et al., 2020; Sankar et al., 2020) to capture the changes across snapshots. However, these approaches neglect fine-grained time information within a snapshot. In contrast, continuous-time methods model graph evolution as a continuous process, capturing all time steps for more precise temporal modeling. These methods often integrate GNNs with specially designed temporal modules, such as temporal random walk (Wang et al., 2021), temporal graph attention (Xu et al., 2020; Rossi et al., 2020), MLP-mixer (Cong et al., 2022) and temporal point processes (Trivedi et al., 2019; Ji et al., 2021; Wen and Fang, 2022). Recently, researchers have proposed a simple and effective architecture called SimpleDyG (Wu et al., 2024), which reformulates dynamic graph modeling as a sequence modeling task. Specifically, it maps the dynamic graph into a series of node sequences and feeds them into a generative sequence model. Subsequently, predicting future events can be framed as a sequence generation problem. However, while these methods provide valuable insights, they are often limited in their ability to adapt to new and evolving patterns in dynamic graphs.
Our work distinguishes itself from prior dynamic graph learning methods through two key innovations. First, while existing approaches predominantly focus on localized temporal contexts or the historical interactions of target nodes, our proposed RAG4DyG employs retrieval-augmented generation mechanisms to retrieve and integrate broader contextual signals from the entire dynamic graph. This approach facilitates a more comprehensive understanding of dynamic interactions, uncovering complex patterns beyond the immediate historical scope of individual nodes. Second, RAG4DyG incorporates a time- and context-aware contrastive learning module for retrieving similar demonstrations, along with a graph fusion strategy to integrate them with the query sequence, enhancing adaptability to new patterns and evolving graph structures.
@@
Retrieval Augmented Generation. Recently, the RAG paradigm has attracted increasing attention (Gao et al., 2023). Specifically, RAG first leverages the retriever to search and extract relevant documents from some databases, which then serve as additional context to enhance the generation process. Related studies have demonstrated the great potential of RAG in various tasks such as language processing (Karpukhin et al., 2020; Jiang et al., 2023), recommendation systems (Ye et al., 2022; Dao et al., 2024), and computer vision (Liu et al., 2023; Kim et al., 2024). In the graph modeling field, existing RAG efforts have primarily focused on static (Mao et al., 2023) and text-attributed graphs to enhance the generation capabilities of Large Language Models (LLMs), supporting graph-related tasks such as code summarization (Liu et al., 2021) and textual graph question answering (He et al., 2024; Hu et al., 2024; Jiang et al., 2024). However, exploiting RAG techniques for dynamic graphs and graphs without textual information remains largely unexplored.
3. Preliminaries
@@
Sequence Mapping of Dynamic Graphs. We denote a dynamic graph as $G=(V,E,F,\mathcal{T})$ comprising a set of nodes $V$, edges $E$, a node feature matrix $F$ if available, and a time domain $\mathcal{T}$. To map a dynamic graph into sequences, we follow SimpleDyG (Wu et al., 2024). Specifically, let $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ denote the set of training samples, where each sample is a pair $(x_{i},y_{i})$, representing the input and output sequences for a target node $v_{i}\in V$. The input $x_{i}$ is a chronologically ordered sequence of nodes that have historically interacted with $v_{i}$, while the output $y_{i}$ is the ground truth interactions that occur following the sequence $x_{i}$. In notations, we have
@@
Problem Formulation. Dynamic graph modeling aims to learn a model that can predict the future interactions of a target node $v_{i}$, given its historical interactions. That is, given $x_{i}$ in Eq. (1), the task is to predict $y_{i}$ in Eq. (2). In our RAG framework, we regard the training samples $D$ as a retrieval pool. Given a target node $v_{q}\in V$, its input sequence $x_{q}$ is referred to as the query sequence. We first retrieve $K$ demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for each query sequence $x_{q}$ based on their contextual and temporal relevance. Next, the retrieved demonstrations $R_{q}$ are used to enrich the input sequence $x_{q}$, which encompasses the historical interactions of the target node $v_{q}$. The augmented input $\{R_{q},x_{q}\}$ is designed to enhance the predictions of future events in $y_{q}$.
4. Proposed Model: RAG4DyG
The RAG4DyG framework enhances dynamic graph modeling by incorporating retrieval-augmented generation techniques to improve predictive accuracy and adaptability. As illustrated in Fig.2, it first adopts SimpleDyG (Wu et al., 2024) to model dynamic graphs as sequences of node interactions, leveraging a Transformer-based model to capture temporal dependencies and predict future interactions (Fig.2(a)). To enrich the modeling process, a time- and context-aware retriever retrieves relevant demonstrations from a retrieval pool $D$ for a given query sequence $x_{q}$. This retriever optimizes two contrastive objectives: a time-aware loss, which employs a temporal decay function $\mu(t_{q},t_{p})$ to prioritize temporally relevant samples, and a context-aware loss, which utilizes sequence augmentation techniques such as masking and cropping to capture structural patterns (Fig.2(b), Sec.4.1). Once the top-$K$ demonstrations are retrieved, they are fused into a summary graph $G_{fus}$, which captures the underlying structural relationships among the retrieved samples. A GNN then processes this graph to generate an enriched representation that is prepended to the query sequence, providing additional context for improved event prediction (Fig.2(c), Sec.4.2). By integrating retrieval and graph fusion, RAG4DyG effectively incorporates temporal and contextual relevance, surpassing existing methods in both transductive and inductive dynamic graph scenarios.
4.1. Time- and Context-Aware Retriever
@@
Retrieval Annotation. To facilitate contrastive training, we automatically annotate the samples in the retrieval pool $D$. For each query sequence $x_{q}$, we annotate its positive sample $x_{p}^{+}$ from the pool $D$ based on their contextual similarity. We leave the detailed annotation process in Sec. 5.1. Specifically, we adopt the sequence model pre-trained in Fig. 2(a) as the encoder and apply mean pooling to obtain sequence representations. Given a query sequence $x_{q}$ and a candidate sequence $x_{p}\in D$, we define their contextual similarity as the dot product of their representations:
(3) $s(x_{q},x_{p})=f(x_{q})^{\top}f(x_{p}),$
where $f(\cdot)$ denotes our encoder.
Time-aware Contrastive Learning. Temporal information reflects the dynamic changes in historical interactions, which is crucial for dynamic graph modeling. We posit that demonstrations closer in time to the query are more relevant than those further away. Consequently, we utilize a time decay function to account for temporal proximity between the query and candidate sequences, as follows.
(4) $\mu(t_{q},t_{p})=\exp(-\lambda|t_{q}-t_{p}|),$
@@
To effectively capture the temporal dynamics of the graph, we incorporate temporal proximity to reweigh the contextual similarity in the contrastive loss:
(5) $h(x_{q},x_{p})$ $=s(x_{q},x_{p})\mu(t_{q},t_{p}).$
@@
(6) $\mathcal{L}_{\text{tcl}}$ $=-\log\frac{\exp(h(x_{q},x_{p}^{+}))/\tau}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp(h(x_{q},x_{j}))/\tau},$
@@
Context-aware Contrastive Learning. To better capture the inherent contextual pattern, we further adopt context-aware contrastive learning with data augmentations. For each sequence, we apply two types of augmentations: masking and cropping, which are widely used for sequence modeling (Devlin et al., 2019; Xie et al., 2022). The masking operator randomly replaces a portion of the tokens in the sequence with a special masking token. The cropping operator randomly deletes a contiguous subsequence from the original sequence, reducing the sequence length while preserving the temporal order of the interactions. These augmentations help the model learn robust representations and capture the inherent structural information of the sequence by focusing on its different parts.
@@
(7) $\mathcal{L}_{\text{ccl}}=-\log\frac{\exp(s(x^{\prime}_{q},x^{\prime\prime}_{q})/\tau)}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp s(x^{\prime}_{q},x^{\prime}_{j})/\tau},$
@@
Training and Inference for Retrieval. The training objective of our retrieval model is defined as:
(8) $\mathcal{L}_{\text{ret}}=\mathcal{L}_{\text{tcl}}+\alpha\mathcal{L}_{\text{ccl}},$
where $\alpha$ is a coefficient that balances between the two losses. During testing, we utilize the updated sequence model to extract sequence representations and perform demonstration ranking based on the contextual similarity between the query and candidates.
4.2. Graph Fusion-based Generator
After the retrieval process, we obtain the top-K demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for the query $x_{q}$. A straightforward approach is to directly concatenate them with the query sequence and input to a sequence generation model for prediction. However, this can lead to a lengthy context that limits the model’s prediction capabilities. More importantly, it neglects the structural patterns among these demonstrations. Thus, we first fuse the demonstrations into a summary graph, process it using a GNN, and then prepend the graph readout from the GNN to the query for subsequent generation.